this is a systems of linear equations in two variables.
How would i go about solving for x and y with the equations ax+y=b and bx-y=a?

Answer by graphmatics(170)

you must solve a 2 by 2 system of equations so first we solve for x using addition of equations
ax+y=b
bx-y=a
___________
(a+b)x = a+b
so dividing both sides by (a+b) we get
x = 1
so a*1 + y = b
y = b-a